OpenStudy (anonymous):

Solve for Y. 3x+7y = 6x+1 If anyone could guide me, that would help so much. Thanks

OpenStudy (anonymous):

Possible answers: y=3x+9 y=7x-1/3 y=3x+1/7 y=9x+1/7

OpenStudy (goldphenoix):

You want 7y by itself. So you want to get rid of 3x. So, (3x + 7y) - 3x = 7y Now, if you subtract 3x on one side of the equation, then you must do it to the other side of the equation. (6x + 1) - 3x = ?

OpenStudy (anonymous):

I got 7y = 3x+1. Is that incorrect?

OpenStudy (goldphenoix):

No. That is correct. :) Alright, now you want just y, not 7y. So what do you think you have to do? Here's a hint: 7y is a coefficient. So another way of writing 7y is 7 * y.

OpenStudy (anonymous):

Would you divide by Y?

OpenStudy (anonymous):

Like, put 7/y

OpenStudy (goldphenoix):

Hmm. Well, 7y / y = 7. Since 7y can be represented as 7 * y. So 7 * y / y = 7. Make sense? You want y, not 7. So what else do you think you have to do?

OpenStudy (goldphenoix):

Our equation looks like, 7y = 3x + 1. If that helps. >_>

OpenStudy (anonymous):

Divide 7y by 7?

OpenStudy (goldphenoix):

Good! Now if you divide 7 on one side of the equation, then you must do it on the other side of the equation. Do you see the answer?

OpenStudy (anonymous):

3x+1/7? :)

OpenStudy (goldphenoix):

Great! I hope this help. :)

OpenStudy (anonymous):

It does! I have one more :)

OpenStudy (anonymous):

Do you think you can help?

OpenStudy (goldphenoix):

Alright. :)

OpenStudy (anonymous):

5(3ax+y)=8ax, -3y for x

OpenStudy (goldphenoix):

So, -3y = x ? Sorry, just making sure.

OpenStudy (anonymous):

I believe so. My paper didn't have a comma after the 8ax, but Ill assume that's what it is.

OpenStudy (anonymous):

There are also some possible answers if this helps: -8y/-23a -8y/-7a -4y/-7a 2y/7a

OpenStudy (goldphenoix):

Let me think for a moment. >_> Gonna check if my answer are correct.

OpenStudy (anonymous):

No problem..

OpenStudy (goldphenoix):

Did the paper say, 5(3ax+y)=8ax -3y for x something like that?

OpenStudy (anonymous):

Yes exactly

OpenStudy (goldphenoix):

Substituting didn't work, I guess that means multiplication. Give me 1m minute.

OpenStudy (anonymous):

Okay Thanks.

OpenStudy (goldphenoix):

I think I know the answer. Its written as 5(3ax+y)=8ax -3y for x Subtract 3y, not comma. >_>

OpenStudy (goldphenoix):

Anyway, distribute. 5(3ax+y) You know how to do the distributive property?

OpenStudy (anonymous):

15ax+5y

OpenStudy (goldphenoix):

Good. Now the equation looks like, 15ax + 5y = 8ax - 3y Combine like terms. Do you know how to do that?

OpenStudy (goldphenoix):

Or simplify. >_>

OpenStudy (anonymous):

7ax+8y?

OpenStudy (goldphenoix):

Hmm. Well, you know that you can subtract 8ax to both side of the equation to get, 7ax + 5y = -3y Now you want 5y on the right side of the equation. Do you know how to do that?

OpenStudy (anonymous):

Ok, ok. subtract 5y to make -8y?

OpenStudy (goldphenoix):

Yep. 7ax = -8y Now you want x. So what do have to do?

OpenStudy (anonymous):

Multiply it by 1?

OpenStudy (goldphenoix):

If you do, then the equation would still be the same. Here's a tip: 7ax can be written as 7 * a * x So your equation can look like: 7 * a *x = -8 * y You want x. So you have to do something about 7 * a .

OpenStudy (anonymous):

put it under 8y?

OpenStudy (goldphenoix):

Hmm. Put what under 8y?

OpenStudy (anonymous):

7a

OpenStudy (goldphenoix):

Great. I think you found your answer.

OpenStudy (anonymous):

Thank you so much. Seriously. You dont know how much I appreciate it. I have to take a College Placement Test and needed help on a few of these, I didn;t understand,

OpenStudy (goldphenoix):

Yeah, no problem. :)